OpenStudy (anonymous):

3x + square root of 6 − 9x = 2

OpenStudy (anonymous):

3x + sqrt(6) - 9x = 2 3x -9x = 2 - sqrt(6) can you do the next step?

OpenStudy (anonymous):

not really.

OpenStudy (anonymous):

its actually 3x + sqrt(6-9x) = 2